OpenStudy (anonymous):

What is the factorization of the polynomial 9x^2 + 6x +1? If it cannot be factored answer n/a

OpenStudy (ybarrap):

Using grouping: $$ 9x^2 + 6x + 1\\ = 9x^2 + 3x + 3x + 1\\ =3x(3x+1)+(3x+1)\\ =(3x+1)(3x+1)\\ =(3x+1)^2\\ $$